# MambaVision: A Hybrid Mamba-Transformer Vision Backbone

### v1

@@ 3.2.2 Layer Architecture
NextViT-small 224 31.7 5.8 3834 82.5
NextViT-B 224    44.8   8.3 2926 83.2
@@
4 Experiments
Image classification experiments are conducted on ImageNet-1K dataset [19]. We followed the standard training recipe of previous efforts [10, 12, 29] to allow for a comparable analysis for performance of different models. Specifically, all models have been trained for $300$ epochs with a cosine decay scheduler which uses an additional $20$ epochs for warmup and cooldown phases, respectively. Furthermore, we use the LAMB optimizer [30] with global batch size of $4096$, an initial learning rate of $0.005$ and weight decay of $0.05$. We note that using the LAMB optimizer resulted in better results compared to traditional AdamW [31], particularly due to its robustness to higher learning rates. We used 32 A100 GPUs for classification tasks.
To evaluate the performance of downstream tasks, we used our pre-trained models as backbones for object detection and instance segmentation and semantic segmentation tasks and used the MS COCO dataset [32] and ADE20K dataset [33], respectively. Specifically, for object detection and instance segmentation, we used Mask-RCNN [34] head with hyperparameters such as ×$\times 3$ LR schedule with an initial learning rate of $0.0001$, a batch size of $16$ and weight decay of $0.05$. For semantic segmentation, we used an UperNet network [35] head with Adam-W [31] optimizer, an initial learning rate of $56e-$ 5 and a global batch size of 16. We used 8 A100 GPUs for all downstream tasks.
Backbone Param (M) FLOPs (G) $\text{AP}^{\text{box}}$ $\text{AP}^{\text{box}}_{50}$ $\text{AP}^{\text{box}}_{75}$ $\text{AP}^{\text{mask}}$ $\text{AP}^{\text{mask}}_{\text{50}}$ $\text{AP}^{\text{mask}}_{\text{75}}$
Mask-RCNN 3$\times$ schedule
Swin-T [10] 48 267 46.0 68.1 50.3 41.6 65.1 44.9
ConvNeXt-T [20] 48 262 46.2 67.9 50.8 41.7 65.0 44.9
MambaVision-T 48 255 46.4 68.3 51.0 41.8 65.4 45.0
Cascade Mask-RCNN 3$\times$ schedule
@@ 4 Experiments
MambaVision-B 145 964 52.8 71.6 57.2 45.7 69.0 49.5
Table 2: Object detection and instance segmentation results with Mask R-CNN and Cascade Mask R-CNN networks using MS COCO dataset [32].
@@ 5.1 Image classification
In Table 1, we present the ImageNet-1K classification results. Specifically, we compare against different families of models such as Conv-based, Transformer-based, Conv-Transformer and Mamba-based and demonstrate that our model outperforms the previous efforts by a large margin, considering ImageNet Top-1 accuracy and image throughput. For instance, when compared to popular models such as ConvNeXt and Swin Transformers, MambaVision-B (84.2%) ConvNeXt-B (83.8%) and Swin-B (83.5%), respectively, while also having significantly better image throughput. We observe similar trends in comparison to Mamba-based models. Specifically, MambaVision-B (84.2%) outperforms VMamba-B (83.9%) despite having considerably higher image throughput. We would also like to note that although our main design goal has been to the accuracy and throughput tradeoff, the MambaVision model variants have much lower FLOPs when compared to similarly-sized counterparts. For instance, MambaVision-B has $56\%$ less GFLOPs than MaxViT-B.
@@ 5.2 Object Detection and Segmentation
We present object detection and instance segmentation results on MS COCO dataset [32] in Table 2. Specifically, we have trained models with different detection sizes to further validate the effectiveness of MambaVision under different scenarios. We also note that our goal is not to achieve SOTA in these tasks but to compare the performance of our backbone with comparably-sized popular vision backbones and validate its effectiveness. With a simple Mask-RCNN detection head, our pretrained MambaVision-T backbone achieves 46.4 and 41.8 in terms of box AP and mask AP, respectively, outperforming both ConvNeXt-T [20] and Swin-T [10] models. With a Cascade Mask-RCNN network, MambaVision-T, MambaVision-S and MambaVision-B outperform the competing counterparts. Specifically, MambaVision models outperform ConvNeXt-T by +0.6 and 0.6,
@@
Table 3: Semantic segmentation results with UperNet [35] networks using ADE20K dataset. For all models, a crop size of $512\times 512$ is used.
ConvNeXt-S by +0.2 and 0.2 and ConvNeXt-B by +0.1 and 0.1 in terms of box AP and mask AP, respectively. Similarly, MambaVision outperforms Swin-T by +0.6 and 0.6, Swin-S by +0.1 and 0.2 and Swin-B by +0.9 and 0.7 in terms of box AP and mask AP, respectively.
We present semantic segmentation benchmarks on the ADE20K dataset [33] in Table 3. For these experiments, we used a UPerNet [35] to be comparable to other models. We observe that MambaVision models outperform similarly-sized competing models for different variants. For instance, MambaVision-T, MambaVision-S and MambaVision-B outperform Swin-T, Swin-S and Swin-B by +0.6, +0.6 and +1.0 in terms of mIoU. Despite the fact that we did not perform extensive optimizations for hyper-parameter tuning of downstream tasks, these results demonstrate the feasibility of MambaVision as a promising backbone for different vision tasks, especially in high-resolution settings.
@@ 5.3 Ablation
In this section, we perform a comprehensive ablation study for a systematic design of MambaVision token mixer. Our goal is to modify the existing Mamba block for computer vision tasks and we evaluate the performance for different tasks such as classification, object detection and instance segmentation and semantic segmentation. All experiments follow a model similar to the architecture layout of the MambaVision-T model as the basis. A shown in Table 4, we start by using the original Mamba formulation with the causal conv layer in the SSM branch (i.e. conv1) and without the additional conv layer in our proposed
@@
conv1 - conv2 - concat 82.3 46.4 41.8 46.6
Table 4: Systematic design of MambaVision token mixer. w/o and concat refer "without" and concatenation. conv1 and conv2 denote the conv operations in the SSM and additional symmetric branch as shown in Fig. 3. The performance is evaluated on ImageNet-1K, MS COCO and ADE20K dataset for classification, object detection and instance segmentation and semantic segmentation, respectively.
symmetric branch (i.e. conv2). As expected, this formulation achieves a suboptimal performance with a Top-1 accuracy of $80.9\%$ ($-1.8\%$), box AP and mask AP of $44.8$ ($-1.6$) and $40.2$ ($-1.6$) and mIoU of $44.2\%$ ($-1.4$). We then replace the causal conv of SSM branch (i.e. conv1) with a regular
Model Top-1
Random 81.3
First $N/2$ layers (SSSSMMMM) 81.5
Mixed layers-1 (SMSMSMSM) 81.4
Mixed layers-2 (MSMSMSMS) 81.6
Last $N/4$ layers (MMMMMMSS) 81.9
Last $N/2$ layers (MMMMSSSS) 82.3
Table 5: Ablation study of on the effectiveness of different hybrid integration patterns using ImageNet-1K dataset. We use the specified patterns for stages 3 and 4. However, only stage 3 (N=8) patterns have been shown in the table for illustration purposes. S and M denote self-attention and MambaVision token mixer blocks, respectively.
conv layer and observe improvements in terms of all metrics due to this change. Furthermore, we add conv2 layer but use the same gating mechanism of Mamba instead of concatenation. This change improves the performance and leads to a Top-1 accuracy of $81.3\%$, box AP and mask AP of $45.3$ and $41.0$ and mIoU of $45.7\%$. Finally, adding concatenation considerably improves the performance in all metrics by $+1.0\%$, $+1.1$, $+0.8$ and $+0.9$ in terms of ImageNet Top-1, box AP and mask AP for MS COCO and mIoU for ADE20K dataset. Hence, this validates our hypothesis that concatenating the outputs from both branches (i.e. SSM and non-SSM) leads to learning richer feature representation and enhancing global context understanding.
Hybrid Pattern
In this section, we comprehensively study the effect of different hybrid integration patterns for self-attention and MambaVision token mixers. For all experiments, the architecture follows MambaVision-T layout with and we keep the models iso-parameter for fair comparison. The patterns are utilized for stages 3 and 4 with hybrid functionality. We begin our study by using a random pattern and achieve a suboptimal Top-1 accuracy of $81.3\%$. This confirms our previous intuition that simply using self-attention without a specific pattern may not be effective. We then utilize the self-attention blocks in the first $N/2$ layers for each stage, with $N$ representing the total number of stage layers, and observe an improvement of $+0.2\%$ ($81.5\%$) in Top-1 accuracy. However, using a mixed layer pattern of self-attention/MambaVision mixers blocks slightly reduces the accuracy by $-0.1\%$ ($81.4\%$). Conversely, reversing the order of mixed layers by using MambaVision/self-attention improves the performance and archives a Top-1 accuracy of ($81.6\%$). We then utilize self-attention blocks in only the last $N/4$ layers of each stage and observe an immediate improvement of $+0.3\%$ ($81.9\%$) in the accuracy. This validates our hypothesis that using self-attention blocks in the final layers of each stage is an effective design. However, its representation learning capability needs to be tuned with respect to MambaVision layers. Increasing the number of self-attention blocks to $N/2$ last layers of each stage achieves the best performance of $82.3\%$.
@@ 6 Conclusion
In this work, we introduced MambaVision which is the first Mamba-Transformer hybrid backbone specifically tailored for vision applications. We proposed re-design of Mamba formulation to enhance global context representation learning capability and presented a comprehensive study of hybrid design integration patterns. MambaVision achieves a new SOTA Pareto front in terms of Top-1 accuracy and image throughput, surpassing Transformer and Mamba-based models by a significant margin. We hope these findings could be the foundation for a new class of hybrid vision models.

### v2

@@ 4 Experiments
Visformer-S [2] 224    40.2    4.8 3676 82.1
NextViT-S [17] 224 31.7 5.8 3834 82.5
NextViT-B [17] 224    44.8   8.3 2926 83.2
NextViT-L [17] 224    57.8 10.8 2360 83.6
EfficientFormer-L1 [18] 224 12.3 1.31 6220 79.2
EfficientFormer-L3 [18] 224    31.4   3.9 2845 82.4
EfficientFormer-L7 [18] 224    82.2 10.2 1359 83.4
@@
Backbone Params (M) FLOPs (G) $\text{AP}^{\text{box}}$ $\text{AP}^{\text{box}}_{50}$ $\text{AP}^{\text{box}}_{75}$ $\text{AP}^{\text{mask}}$ $\text{AP}^{\text{mask}}_{\text{50}}$ $\text{AP}^{\text{mask}}_{\text{75}}$
@@
MambaVision-B 145 964 52.8 71.3 57.2 45.7 68.7 49.4
Table 2: Object detection and instance segmentation benchmarks using Cascade Mask R-CNN [13] on MS COCO dataset [19]. All models are trained by using a 3×3$\times$3 × schedule and a crop resolution of $1280\times 800$.
@@ 5.1 Image classification
In Table 1, we present the ImageNet-1K classification results. Specifically, we compare against different families of models such as Conv-based, Transformer-based, Conv-Transformer, and Mamba-based architectures and demonstrate that our model outperforms the previous efforts by a large margin, considering ImageNet Top-1 accuracy and image throughput. For example, MambaVision-B achieves higher accuracy (84.2%) compared to ConvNeXt-B (83.8%) and Swin-B (83.5%), while also having significantly better image throughput. We observe similar trends in comparison to Mamba-based models. Specifically, MambaVision-B (84.2%) outperforms VMamba-B (83.9%) despite having considerably higher image throughput. We observe similar trends in performance comparisons with respect to other Mamba-based models. In addition, we would also like to note that although our main design goal has been to optimize the accuracy and throughput tradeoff, the MambaVision model variants have much lower FLOPs when compared to similarly-sized counterparts. For instance, MambaVision-B has 56% less GFLOPs than MaxViT-B.
@@ 5.2 Object Detection and Segmentation
We evaluate our model’s object detection and instance segmentation performance on the MS COCO dataset [19], as shown in Table 2. To comprehensively validate MambaVision’s effectiveness, we trained models of varying sizes and compared them with popular vision backbones of comparable scale under identical conditions. Using a Cascade Mask R-CNN [13] head, all variants of MambaVision demonstrated superior performance compared to their counterparts. Specifically, MambaVision models outperform ConvNeXt-T by +0.7 and +0.6, ConvNeXt-S by +0.4 and +0.2 and ConvNeXt-B by +0.1 and +0.1 in terms of box Average Precision (AP) and mask AP, respectively. Similarly, MambaVision outperforms Swin-T by +0.7 and +0.6, Swin-S by +0.4 and +0.2 and Swin-B by +0.9 and +0.7 in terms of box AP and mask AP, respectively. For semantic segmentation, we evaluated the performance on the ADE20K dataset [39] using UPerNet [34], as shown in Table 3. We observe that MambaVision models outperform similarly-sized competing models for different variants. For instance, MambaVision-T, MambaVision-S, and MambaVision-B outperform Swin-T, Swin-S, and Swin-B by +1.5, +0.6, and +1.0 in terms of mIoU, respectively. Notably, these improvements were achieved without extensive hyperparameter optimization for downstream tasks, highlighting MambaVision’s potential as a robust backbone for various vision tasks, particularly in high-resolution scenarios. Moreover, our approach consistently attains higher mIoU than Focal Transformers across all scales while having comparable model sizes.
@@
Table 3: Semantic segmentation results with UperNet [34] model using ADE20K dataset. All models are trained using a crop resolution of $512\times 512$.
@@ 5.3 Ablation
Large-scale Training on ImageNet-21K
For the first time in any Mamba-based approach, our work (MambaVision) has scaled training to the large ImageNet-21K dataset with significantly bigger model sizes. As demonstrated in Fig. 4 the results are promising. Specifically, we observe meaningful improvements for the smaller MambaVision-B model (97.7M parameters), whose Top-1 accuracy increases from 84.2% to 84.9% at 224 resolution. In addition, pre-training and fine-tuning MambaVision-L raises its Top-1 accuracy from 85% to 86.1% at 224 resolution. We have also introduced a larger variant, MambaVision-L3 (739.6M parameters), which attains Top-1 accuracies of 87.3% and 88.1% at 256 and 512 resolutions, respectively. These results validate the scalability of our model across larger datasets, varying model sizes, and different image resolutions. To the best of our knowledge, MambaVision represents the first successful scaling of a Mamba-based vision architecture to ImageNet-21K with strong performance. This ability to scale is critical for real-world scenarios that rely on massive datasets, where bigger, more capable models are needed to achieve robust performance. We anticipate that MambaVision’s proven scalability will further encourage the adoption of Mamba-based models in industrial and large-scale research applications.
@@
We conducted a comprehensive ablation study to systematically design the MambaVision token mixer. Our investigation focused on adapting the Mamba block for computer vision tasks, evaluating performance across classification, object detection, instance segmentation, and semantic segmentation. All experiments used a model architecture based on MambaVision-T configuration. As shown in Table 4, we began with the original Mamba formulation, which includes a causal convolution layer in the SSM branch (conv1) but lacks the additional convolution layer in our proposed symmetric branch (conv2). This baseline configuration achieved suboptimal performance across all metrics, with ImageNet Top-1 accuracy of 80.9% (-1.8%), MS COCO box AP of 44.8 (-1.6) and mask AP of 40.2 (-1.6), and ADE20K mIoU of 44.2% (-1.4). We then replaced the causal convolution in the SSM branch (conv1) with a regular convolution layer, which improved performance across all metrics. Subsequently, we added conv2 layer while maintaining Mamba’s original gating mechanism instead of concatenation, resulting in ImageNet Top-1 accuracy of 81.3%, MS COCO box AP of 45.3 and mask AP of 41.0, and ADE20K mIoU of 45.7%. Finally, implementing concatenation led to substantial improvements across all metrics, with gains of +1.0% in ImageNet Top-1, +1.1 in box AP and +0.8 in mask AP for MS COCO, and +0.9 in mIoU for ADE20K. These results validate our hypothesis that concatenating outputs from both branches (SSM and non-SSM) enables the model to learn richer feature representations and enhance global context understanding.
Figure 4: Performance scalability of MambaVision ImageNet-21K pretrained models with varying model sizes and resolutions.
Hybrid Pattern
We conducted a comprehensive study examining various hybrid integration patterns between self-attention and MambaVision token mixers. All experiments maintained the MambaVision-T architecture layout with iso-parameter models for fair comparison, implementing hybrid functionality in stages 3 and 4. Initial experiments with a random integration pattern yielded suboptimal results with a Top-1 accuracy of 81.3%, confirming our intuition that arbitrary self-attention placement may be ineffective. When we positioned self-attention blocks in the first N/2 layers of each stage (where N represents the total number of stage layers), performance improved by +0.2% (81.5%). A mixed layer pattern alternating between self-attention and MambaVision mixer blocks showed a slight performance decrease of -0.1% (81.4%), while reversing this order to MambaVision/self-attention improved accuracy to 81.6%. Placing self-attention blocks in only the last N/4 layers of each stage yielded a significant improvement of +0.3% (81.9%), supporting our hypothesis that self-attention is most effective in the final layers. Further optimization revealed that extending self-attention to the last N/2 layers of each stage achieved the best performance at 82.3%, indicating the importance of carefully balancing self-attention blocks with MambaVision layers for optimal representation learning.
Table 4: Systematic design of MambaVision token mixer. w/o and concat refer to "without" and concatenation. Conv1 and conv2 denote the conv operations in the SSM and additional symmetric branch as shown in Fig. 3. COCO experiments are performed using Mask-RCNN [13] head and ×1 LR schedule.
@@
conv1 - conv2 - concat 82.3 46.4 41.8 46.0
Figure 5: Visualizations of MambaVision’s self-attention layers showing how the model learns to focus on semantically meaningful regions via attention maps (middle) and overlays (right).
Table 5: Ablation study of on the effectiveness of different hybrid integration patterns. S and M denote self-attention and MambaVision token mixer blocks, respectively.
Model Pattern Params (M) Top-1
Random - 31.8 81.3
First $N/2$ layers SSSSMMMM 31.8 81.5
Mixed layers-1 SMSMSMSM 31.8 81.4
Mixed layers-2 MSMSMSMS 31.8 81.6
Last $N/4$ layers MMMMMMSS 31.8 81.9
Last $N/2$ layers MMMMSSSS 31.8 82.3
Interpretability
To better understand how MambaVision processes visual information, we visualize the attention maps from the self-attention layers in the final stages. As shown in Fig. 5, these visualizations reveal that the model learns to focus on semantically meaningful regions without explicit supervision. In the aircraft example, the attention clearly highlights the entire plane body, suggesting effective capture of object boundaries. For the bird image, we observe concentrated attention on distinctive features like the head and tail regions, demonstrating the model’s ability to identify fine-grained details. In the case of object-human interaction (bottom row), the attention map shows strong activation on both the subject and the object being held, indicating that the self-attention layers successfully model relationships between different elements in the scene. These visualizations support our architectural design choice of using self-attention blocks in the final stages to capture global context and long-range dependencies.
@@ 6 Conclusion
In this work, we introduced MambaVision which is the first Mamba-Transformer hybrid backbone specifically tailored for vision applications. We proposed re-design of Mamba formulation to enhance global context representation learning capability. MambaVision achieves a new SOTA Pareto front in terms of Top-1 accuracy and image throughput, outperforming Transformer and Mamba-based models by a significant margin. Through extensive experimentation across multiple vision tasks, including classification, detection and segmentation, we demonstrated the versatility and effectiveness of our approach. Our systematic analysis of integration patterns revealed that positioning self-attention blocks in the final layers significantly improves the model’s ability to capture long-range dependencies while maintaining efficiency. Furthermore, we successfully scaled MambaVision to ImageNet-21K pretraining, achieving strong performance that matches SOTA models, demonstrating its potential for large-scale vision applications. The success of MambaVision in addressing the limitations of pure Mamba-based architectures while leveraging their strengths opens new possibilities for vision backbone design. We hope these findings could be the foundation for a new class of hybrid vision models.
Appendix
Figure S.1: Visualization of MambaVision’s attention patterns. Each row contains two example cases, with each case showing a triplet of: (left) original input image, (middle) attention heat map, and (right) attention overlay on the input image. The examples showcase diverse scenarios: containers and spiders (row 1), aircraft and birds (row 2), marine life and snakes (row 3), groomed dogs and extreme sports (row 4), poultry and snakes (row 5), and arachnids and outdoor activities (row 6). The attention maps reveal how MambaVision effectively localizes key semantic regions and object boundaries across this wide range of categories.
G Ablation Study
To determine the optimal window size for MambaVision models, we study its impact on the performance of MambaVision-T in different tasks such as image classification, object detection and instance segmentation. Given $Q,K,V$ as the query, key and value tensors respectively, self-attention is computed according to
$d_{h}$ represents the number of attention heads. If the input size is larger than the window size, the attention is computed in the local windows. Specifically, we study two different architectures with window sizes of 7 and 14 in their stage 3 of the model. We also measure image throughput for the task of image classification with a batch size of 128. As presented in Table S.1, our analysis reveals that increasing the window size to 14 offers a favorable trade-off between performance and computational cost. While maintaining nearly identical throughput (6298 img/s vs. 6318 img/s), the larger window size achieves consistent improvements across all vision benchmarks: ImageNet top-1 accuracy increases to 82.3%, COCO mask AP improves to 41.8%. These gains, though modest, come with minimal computational overhead on modern hardware such as the NVIDIA A100 GPU. Based on this empirical evidence, we selected 14 and 7 as our default window sizes, as this combination provides better vision understanding capabilities while preserving the model’s efficiency. The negligible 0.3% decrease in throughput is well justified by the improved performance in various vision tasks.
  Model Window Size Throughput (img/s) ImageNet COCO
[…]
MambaVision-T 7,7 6318 82.2 46.4 41.7
MambaVision-T 14,7 6298 82.3 46.4 41.8
Table S.1: Ablation study on window size for MambaVision model’s performance. Experiments on COCO dataset [19] are performed using Mask-RCNN [13] head and ×1 LR schedule. Throughput is measured for image classification on a single NVIDIA A100 GPU with batch size 128.
H Architecture Details
In Table S.2, we present the comprehensive architectural specifications of MambaVision variants. The backbone follows a hierarchical design with 4 stages, each employing convolutional down-sampling operations that progressively reduce spatial resolution by a factor of two. A key innovation in our architecture appears in Stages 3 and 4, where we introduce a hybrid design that synergistically combines Mamba-based sequence modeling with self-attention mechanisms. This hybrid approach leverages Mamba’s efficient sequence processing capabilities while benefiting from the global context modeling strengths of self-attention layers. Each variant (T, S, B, and L) maintains this fundamental structure while scaling the channel dimensions and layer counts to achieve different complexity-performance trade-offs.
Output Size
(Downs. Rate)
MambaVision-T MambaVision-S MambaVision-B MambaVision-L
Stem 112$\times$112 (2$\times$) $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:32, S:2}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:64, S:2}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:64, S:2}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:64, S:2}\end{bmatrix}$$\times$ 1
$\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:80}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:96}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:128}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:196}\end{bmatrix}$$\times$ 1
Stage 1 56$\times$56 (4$\times$) Conv, C:160, S:2 Conv, C:192, S:2 Conv, C:256, S:2 Conv, C:392, S:2
$\begin{bmatrix}\text{ResBlock}\\ \text{C:160}\end{bmatrix}$$\times$1, $\begin{bmatrix}\text{ResBlock}\\ \text{C:192}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:256}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:392}\end{bmatrix}$$\times$ 3,
Stage 2 28$\times$28 (8$\times$) Conv, C:320, S:2 Conv, C:384, S:2 Conv, C:512, S:2 Conv, C:768, S:2
$\begin{bmatrix}\text{ResBlock}\\ \text{C:320}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:384}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:512}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:768}\end{bmatrix}$$\times$ 3,
Stage 3 14$\times$14 (16$\times$) Conv, C:640, S:2 Conv, C:768, S:2 Conv, C:1024, S:2 Conv, C:1568, S:2
$\begin{bmatrix}\text{MV}\\ \text{C:640}\end{bmatrix}$$\times$ 4, $\begin{bmatrix}\text{SA}\\ \text{C:640, head:8}\end{bmatrix}$$\times$ 4 , $\begin{bmatrix}\text{MV}\\ \text{C:768}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:768, head:8}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{MV}\\ \text{C:1024}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:1024, head:8}\end{bmatrix}$$\times$ 4, $\begin{bmatrix}\text{MV}\\ \text{C:1568}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:1568, head:16}\end{bmatrix}$$\times$ 3,
Stage 4 7$\times$7 (32$\times$) Conv, C:1280, S:2 Conv, C:1536, S:2 Conv, C:2048, S:2 Conv, C:3136, S:2
$\begin{bmatrix}\text{MV}\\ \text{C:1280}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:1280, head:16}\end{bmatrix}$$\times$ 4, $\begin{bmatrix}\text{MV}\\ \text{C:1536}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:1536, head:16}\end{bmatrix}$$\times$ 2, $\begin{bmatrix}\text{MV}\\ \text{C:2048}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:2048, head:16}\end{bmatrix}$$\times$ 2, $\begin{bmatrix}\text{MV}\\ \text{C:3136}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:3136, head:32}\end{bmatrix}$$\times$ 2,
Table S.2: Architecture configurations of MambaVision models. SA and MV refer to self-attention and MambaVision mixer blocks respectively. BN denote Batch Normalization.
I Training Details
Image classification experiments are conducted on the ImageNet-1K dataset [4]. All models have been trained for 300 epochs using 32 A100 GPUs, with LAMB optimizer, batch size of 4096, and learning rate of 4e-3. The self-attention formulation in stages 3 and 4 of all MambaVision variants use a window size of 14 and 7, respectively. To evaluate the performance of downstream tasks, we used our pre-trained models as backbones for object detection, instance segmentation, and semantic segmentation tasks using the MS COCO dataset [19] and ADE20K dataset [39], respectively. For all downstream tasks, we used an AdamW optimizer and batch size of 16. Specifically, for object detection and instance segmentation, we used the Cascade Mask-RCNN [13] head with hyperparameters such as ×3 LR schedule. For semantic segmentation, we used a UperNet network [34] segmentation head.
J Interpretability
To demonstrate the interpretability of MambaVision models, we visualize the attention patterns learned by our model across diverse object categories. Figure S.1 presents a comprehensive analysis of attention mechanisms through paired examples.
Our paired visualization analysis reveals several key insights about MambaVision’s visual processing capabilities:
  • Consistent Pattern Recognition: Each triplet (input-heatmap-overlay) demonstrates how the model maintains consistent attention patterns across different instances of similar object categories.
  • Contextual Understanding: The paired examples within each row often represent contrasting scenarios (e.g., man-made objects vs. natural subjects), showing the model’s adaptability across domains.
  • Fine-grained Detail: The attention heat maps precisely highlight discriminative features, from the texture of animal fur to the structural elements of vehicles and containers.
  • Robust Localization: Across all example pairs, the overlaid visualizations demonstrate accurate object boundary detection, regardless of the subject’s position or background complexity.